# SurgRIPE challenge: Benchmark of Surgical Robot Instrument Pose Estimation

### v1

@@ 7 Analysis of the results
In the first category, TUDU and MVL_3S aim to estimate the 3D translation and rotation of the target object directly from image features in a single forward pass. They use segmentation to generate a 2D bounding box for 2D localisation. The direct estimation pipeline offers simplicity but often struggles with ambiguities, complex rotations, and occlusions due to the absence of intermediate geometric reasoning. These limitations cause the higher rotation error and performance drop under occlusion, which is reflected by the low ADD and Ave Acc metrics demonstrated in results of the previous section. As shown in the 6DoF trajectories in Fig. 11 -  14, MVL_3S achieved a larger average rotation error than IGTUM and ICL.
Two-stage methods, like ICL, EUT and PVNet, involve extracting intermediate representations, such as keypoints or segmentation masks, before the final pose estimation. Instead of using a deep learning model to estimate the pose directly, these models use intermediate results and the PnP solver Lepetit et al. (2009) to compute 2D-3D keypoint correspondences which are processed further for 6DoF pose estimation. The two-stage architecture makes these methods more robust to occlusions and complex geometries. However, errors in the intermediate results like incorrect localisation of a single keypoint in the first stage, can propagate to the second stage in the PnP solver, resulting in a significant overall error. This property is also reflected in their high Average Accuracy and high translation and rotation error in the results section. According to the estimated trajectories shown for the Z-axis translation estimation in Fig. 11-14, two-stage methods like ICL generate more frequent outliers instead of consistent trajectories compared to the direct prediction methods.
@@
Figure 11: Trajectories comparison in LND Test Without Occlusion
Figure 12: Trajectories comparison in MBF Test Without Occlusion
Figure 13: Trajectories comparison in LND Test With Occlusion
Figure 14: Trajectories comparison in MBF Test With Occlusion
@@ 8 Discussion
8.1 Dataset
8.1.1 Ground truth data accuracy
To evaluate the accuracy of our ground truth pose generation, we created simulated images with known instrument pose information. More specifically, we used the VisionBlender simulation platform Cartucho et al. (2020) to generate images of the instrument with a keydot marker, accompanied by their actual pose. As shown in Fig. 15, the keydot marker moved along with the tool 3D model. To guarantee consistency between the simulation images and real endoscopic images, the same camera intrinsic matrix was used as in the LND subset. To generate ground truth tool pose estimation for the simulated data, we estimated the pose of the keydot marker by analysing the simulation images and followed the procedure explained in Section 2.2. Among the 50 samples of simulated data, the translation error between the estimated and actual instrument pose is 0.253 mm and the rotation error is 0.302 degrees. This verifies the accuracy of our generated ground truth data.
Figure 15: Simulation sample images
@@ 8.2 Future prospects
Dataset Improvement: SurgRIPE dataset focuses on instrument pose estimation for monocular endoscopic images. Considering da Vinci®endoscope can capture stereo images, more stereo images could be collected in the future to utilize extra depth information from stereo images. In addition, the dataset can be improved by adding more realistic tissue as background to simulate real surgery.
@@ Acknowledgments
Funding The authors are grateful for the sponsorship from Intuitive® and for the prize awards.

### v2

@@ 8 Analysis of the results
In the first category, TUDU and MVL_3S aim to estimate the 3D translation and rotation of the target object directly from image features in a single forward pass. They use segmentation to generate a 2D bounding box for 2D localisation. The direct estimation pipeline offers simplicity but often struggles with ambiguities, complex rotations, and occlusions due to the absence of intermediate geometric reasoning. These limitations cause the higher rotation error and performance drop under occlusion, which is reflected by the low ADD and Avg Acc metrics demonstrated in results of the previous section. As shown in the 6DoF trajectories in Fig. 11-14, MVL_3S achieved a larger average rotation error than IGTUM and ICL.
Two-stage methods, like ICL, EUT and PVNet, involve extracting intermediate representations, such as keypoints or segmentation masks, before the final pose estimation. Instead of using a deep learning model to estimate the pose directly, these models use intermediate results and the PnP solver (Lepetit et al., 2009) to compute 2D-3D keypoint correspondences which are processed further for 6DoF pose estimation. The two-stage architecture makes these methods more robust to occlusions and complex geometries. However, errors in the intermediate results like incorrect localisation of a single keypoint in the first stage, can propagate to the second stage in the PnP solver, resulting in a significant overall error. This property is also reflected in their high Average Accuracy and high translation and rotation error in the results section. According to the estimated trajectories shown for the Z-axis translation estimation in Fig. 11-14, two-stage methods like ICL generate more frequent outliers instead of consistent trajectories compared to the direct prediction methods.
@@
Figure 11: Trajectory comparison in LND Test Without Occlusion.
Figure 12: Trajectory comparison in MBF Test Without Occlusion.
Figure 13: Trajectory comparison in LND Test With Occlusion.
Figure 14: Trajectory comparison in MBF Test With Occlusion.
@@ 9 Discussion
9.1 Role of Vision-based Deep Learning in Surgical Instrument Pose Estimation
The participants’ results demonstrate the promising potential of applying advanced deep learning technology in surgical instrument pose estimation, especially in the occluded scenarios which are common but challenging in real surgery. Compared with the traditional tracking methods relying on extra hardware and markers, the vision-based methods provide more flexibility and cost-efficiency. However, there is still clear scope for improvement in further application. Although some two-stage methods could provide intermediate results, deep learning methods still lack interpretability and uncertainty estimation. In addition, all the participants’ methods require retraining for every new instrument, while the ground truth 6DoF pose annotation is quite time-consuming and limited by the instrument size. In the future, generalizable pre-trained model could be implemented to solve the issue.
9.2 Ground truth data accuracy and error analysis
To evaluate the accuracy of our ground truth pose generation, we created simulated images with known instrument pose information. More specifically, we used the VisionBlender simulation platform (Cartucho et al., 2020) to generate images of the instrument with a keydot marker, accompanied by their actual pose. As shown in Fig. 15, the keydot marker moved along with the tool 3D model. To guarantee consistency between the simulation images and real endoscopic images, the same camera intrinsic matrix was used as in the LND subset. To generate ground truth tool pose estimation for the simulated data, we estimated the pose of the keydot marker by analysing the simulation images and followed the procedure explained in Section 2.2. Among the 50 samples of simulated data, the translation error between the estimated and actual instrument pose is 0.253 mm and the rotation error is 0.302 degrees. This verifies the accuracy of our generated ground truth data.
Figure 15: Simulation sample images.
9.3 Background Scene Simulation
The background scene was simulated with a realistic high-fidelity abdominal phantom in our experiments. To create diversity of the background conditions, video data was captured from different areas of the phantom corresponding to different organs, using the Da Vinci Si camera. Varying lighting conditions were generated by adjusting the light source intensity on the Da Vinci between 40% and 100% with a step of 10%. In addition, a ceiling light source was used to illuminate our scene. We adjusted the intensity and the wavelength of the light source, and we also moved it to different locations above our experimental set up to further increase the variability of the lighting conditions in our captured video data.
9.4 Monocular Image Capture
In this work, we chose to capture only the left camera images for data collection due to practical and methodological reasons. First, the physical marker used for 6DoF pose ground truth is only reliably detected within limited viewing angles, and requiring visibility in both stereo views would constrain pose diversity. Second, applying image inpainting to both stereo images could introduce inconsistencies that affect stereo matching and depth estimation. Finally, using stereo data increases the risk of calibration errors between cameras, which could degrade data quality. Focusing on the left images avoids these issues and improves dataset robustness.
@@ 9.5 Future prospects
Dataset Improvement: SurgRIPE dataset focuses on instrument pose estimation for monocular endoscopic images. Considering da Vinci®endoscope can capture stereo images, more stereo images could be collected in the future to utilize extra depth information from stereo images. In addition, the dataset can be improved by adding more realistic tissue as background to simulate real surgery. So far, the target for the pose estimation is the wrist joint between the shaft and the end effector of the tool. The tool pose is estimated with respect to the camera coordinate system, which is our reference coordinate system. Since the end effector can only rotate along the z-axis of the wrist joint, the pose of the tool tip can be easily recovered using the estimated wrist joint pose by detecting the 2D position of the tool tip and applying simple 3D geometry.
@@ 10 Conclusions
10.1 CRediT authorship contribution statement
Haozheng Xu: Conceptualization, Methodology, Software, Resources, Data Curation, Formal analysis, Investigation, Validation, Visualization, Funding acquisition, Writing - Review & Editing, Project administration. Alistair Weld: Software, Formal analysis, Presentation, Investigation, Validation, Visualization, Writing - Review & Editing, Project administration. Chi Xu: Software, Validation, Presentation, Visualization. Alfie Roddan: Software, Validation, Presentation, Visualization. João Cartucho: Writing - Original Draft. Mert Asim Karaoglu: Investigation, Writing - Original Draft. Alexander Ladikos: Investigation, Writing - Original Draft. Yangke Li: Investigation, Writing - Original Draft. Yiping Li: Investigation, Writing - Original Draft. Daiyun Shen: Investigation, Writing - Original Draft. Shoujie Yang: Investigation, Writing - Original Draft. Geonhee Lee: Investigation, Writing - Original Draft. Seyeon Park: Investigation, Writing - Original Draft. Jongho Shin: Investigation, Writing - Original Draft. Young-Gon Kim: Investigation, Writing - Original Draft. Lucy Fothergill: Investigation, Writing - Original Draft. Dominic Jones: Investigation, Writing - Original Draft. Pietro Valdastri: Investigation, Writing - Original Draft. Duygu Sarikaya: Investigation, Writing - Original Draft. Stamatia Giannarou: Conceptualization, Writing - Review & Editing, Supervision.
@@ Acknowledgments
Funding The authors are grateful for the sponsorship from Intuitive® for the prize awards.